OpenStudy (anonymous):

The perimeter of a rectangle is 58 meters. The width is 6 less than the length. Which system of equations can you use to find the width of the rectangle? A. lw= 58 l - 2=6 B. 2(l + w)=58 6 + 2w= 2l C. 2l + 2w= 58 -l + w= -6 D. 2(l + 2)=58 lw=6

OpenStudy (anonymous):

if l is the length and w is the width; w+6 = l

OpenStudy (texaschic101):

perimeter of rectangle : P = 2L + 2W P = 58 W = L - 6 .....-L + W = -6 I am going with C